2. Arrange the same data above in a simple frequency distribution table with X values listed in 7 to 15 categories. You decide what would make the most reasonable display based on the criteria we discussed in class. (Hint: using an arrangement like your stem and leaf display groupings would not be a bad idea.) Leave enough room to display additional values as described in #4.46 below. 3. Determine N for the above data. N- 4. Indicate the relative frequency for each interval category of X on the table you created in #2. 5. Indicate the cumulative frequency for each category of X. 6. Compute the relative cumulative frequencies for each X interval.40-49 30-39 25 20-29 100 80 40 10-19 0-9 n-250
Answer #1

N for data, N = 250

X f    Relative frequency Cumulative frequency Relative Cumulative frequency
40-49 5 = 5/250 = 0.02 = 5 = 0.02
30-39 25 = 25/250 = 0.1 = 5 + 25 = 30 = 0.02 + 0.10 = 0.12
20-29 100 = 100/250 = 0.4 = 30 + 100 = 130 = 0.12 + 0.04 = 0.52
10-19 80 = 80/250 = 0.32 = 130 + 80 = 210 = 0.52 + 0.32 = 0.84
0-9 40 = 40/250 = 0.16 = 210 + 40 = 250 = 0.84 + 0.16 = 1.00